# Minchul Kim, Anil Jain, Xiaoming Liu

### v2

@@ 4.5 Feature Fusion in Face Recognition
In template-based face recognition, multiple face images of the same identity—often captured under varying conditions of pose, illumination, resolution, and occlusion—must be fused into a single, compact representation to enable efficient and accurate comparison. This fusion scenario commonly arises in gallery settings, where multiple still images (or media) per subject must be aggregated into a unified template.
@@
A second, and increasingly popular, scenario involves video-based face recognition, where frames extracted from probe video sequences are fused into a single representation. This use case poses unique challenges, as it often requires online (on-the-fly) feature fusion to support real-time applications such as surveillance or mobile authentication. Despite differing temporal constraints, both still-image and video-based fusion share the core objective: to generate robust and compact representations that preserve discriminative identity cues.
@@
Feature fusion is a critical step in this process, as it determines how the information from diverse images of the same person is aggregated into a unified descriptor. Naive methods like average or max pooling treat all feature embeddings equally, which can dilute discriminative cues by giving equal importance to low-quality or redundant images. Therefore, effective feature fusion must not only compress but also intelligently filter, weight, and adapt to the content of the input set. The ability to generate robust, order-invariant, and compact template representations directly impacts FR performance, especially in unconstrained or real-time scenarios.
@@
Notably, one of the earliest contributions to video-based face recognition uses adaptive hidden Markov models to model temporal dynamics and perform recognition based on entire video sequences [72]. Over the years, feature fusion in FR has evolved from simple averaging techniques to highly adaptive and context-aware neural aggregation frameworks. Early methods such as the Neural Aggregation Network (NAN) [139] establishes the value of learning quality-aware weights through attention, enabling robust, order-invariant representations for face templates. This sets the stage for more sophisticated models like Multicolumn Networks [140] and C-FAN [141], which incorporate fine-grained quality analysis, either by separating visual and contextual importance or by estimating weights for individual feature channels. These advances are instrumental in improving FR accuracy on challenging template-based datasets such as IJB-C [24].
@@
Recent efforts have shifted toward scaling and generalizing feature fusion across different operating conditions. Methods such as CAFace [71], CoNAN [142] and ProxyFusion [143] focus on maintaining performance even when templates contain hundreds or thousands of diverse images. Meanwhile, practical methods like Norm Pooling [144] demonstrate that even simple heuristics can be powerful in multi-domain scenarios, especially when training data is limited. A consistent trend across these innovations is the move toward scalable and fast strategies that can tailor to long videos, achieving robustness without sacrificing scalability.
@@ 5 State of the art in Face Recognition
The evaluation of face recognition (FR) systems broadly falls into three categories:
Benchmark Evaluation: Self-evaluation by algorithm developers on public benchmark datasets such as LFW, IJB-C, and TinyFace. Some examples are shown in Fig. 14.
Technology Evaluation: Third-party independent testing conducted by organizations like NIST on operational datasets (e.g., mugshots, visa photos, border kiosk captures) that are not publicly available. Developers submit their algorithms for unbiased evaluation.
Scenario Evaluation: Deployment of FR systems in real-world operational settings, such as airport boarding gates, where practical concerns such as device placement, lighting, usability, and capture quality become critical alongside algorithmic performance.
In this section, we present two types of evaluations: (1) Benchmark evaluations and (2) Technology evaluations conducted by NIST.
TABLE IV: Performance on CFP-FP [27] Dataset
@@
GFace [145] IResNet-50 GCE (LO) Casia-WebFace 97.44
CosFace [19] ResNet100 CosFace [19] MS1MV2 98.13
ArcFace [20] ResNet101 ArcFace MS1MV2 98.27
MV-Softmax [146] ResNet100 MV-Softmax MS1MV2 98.28
CurricularFace [21] ResNet101 CurricularFace MS1MV2 98.37
TransFace-B [147] ResNet100 ArcFace MS1MV2 98.39
MagFace [63] ResNet100 MagFace MS1MV2 98.46
AdaFace [22] ResNet101 AdaFace MS1MV2 98.49
CQA-Face [148] ResNet100 CQA-Face MS1MV2 98.49
UniFace [65] ResNet100 UniFace [65] MS1MV2 98.63
URL [149] ResNet101 URL MS1MV2 98.64
LGAF [150] ResNet100 ArcFace MS1MV2 98.77
ArcFace [20] ResNet50 ArcFace Glint360K 98.77
ViT-S [102] ViT-S ArcFace Glint360K 98.85
@@
TransFace-S [147] ViT-S ArcFace Glint360K 98.91
AdaFace [22] ViT AdaFace WebFace4M 98.94
KP-RPE [39] ViT AdaFace WebFace4M 99.01
ViT-B [102] ViT-B ArcFace Glint360K 99.02
AdaFace [22] ResNet101 AdaFace MS1MV3 99.03
@@
AdaFace [22] ViT AdaFace MS1MV3 99.06
ArcFace [20] ResNet101 ArcFace WebFace4M 99.06
KP-RPE [39] ViT ArcFace WebFace4M 99.09
@@
KP-RP [39]E ViT AdaFace MS1MV3 99.11
GFace [145] IResNet-100 GCE (LO) MS1MV3 99.12
@@
AdaFace [22] ResNet101 AdaFace WebFace4M 99.17
TransFace-B [147] ViT-B ArcFace Glint360K 99.17
AdaFace [22] ResNet101 AdaFace WebFace12M 99.24
KP-RPE [39] ViT AdaFace WebFace12M 99.30
TransFace-L [147] ViT-L ArcFace Glint360K 99.32
TABLE V: Performance on IJB-C [24] Dataset
@@
ArcFace [20] ResNet101 ArcFace [20] MS1MV2 96.03
MagFace [63] ResNet101 MagFace [63] MS1MV2 95.81
@@
CurricularFace [21] ResNet101 CurricularFace MS1MV2 96.10
@@
TransFace-S [147] ViT-S ArcFace MS1MV2 96.45
TransFace-B [147] ViT-B ArcFace MS1MV2 96.55
TransFace-L [147] ViT-L ArcFace MS1MV2 96.59
@@
ARoFace [151] ResNet101 ArcFace MS1MV2 96.66
ElasticFace [64] ResNet101 ElasticFace MS1MV2 96.65
TopoFR [152] ResNet101 TopoFR MS1MV2 96.95
GFace [145] ResNet101 TopoFR MS1MV2 96.96
AdaFace [22] ResNet101 AdaFace MS1MV2 97.09
KP-RPE [39] ViT-B CosFace WebFace4M 96.98
TopoFR [152] ResNet200 TopoFR MS1MV2 97.08
AdaFace [22] ViT-B AdaFace MS1MV3 97.10
KP-RPE [39] ViT-B AdaFace WebFace4M 97.13
AdaFace [22] ViT-B AdaFace WebFace4M 97.14
KP-RPE [39] ViT-B AdaFace MS1MV3 97.16
KP-RPE [39] ViT-B ArcFace WebFace4M 97.21
PartialFC [79] ResNet101 ArcFace WebFace4M 97.22
CatFace [153] ResNet101 CatFace MS1MV2 97.43
AdaFace [22] ResNet101 AdaFace WebFace4M 97.39
ARoFace [151] ResNet101 AdaFace WebFace4M 97.51
AdaFace [22] ResNet101 AdaFace WebFace12M 97.66
PartialFC [79] ResNet101 ArcFace WebFace12M 97.58
ARoFace [151] ResNet101 AdaFace WebFace12M 97.60
TopoFR [152] ResNet101 TopoFR Glint360K 97.60
KP-RPE [39] ViT-B AdaFace WebFace12M 97.82
PartialFC [79] ResNet101 ArcFace WebFace42M 97.82
TopoFR [152] ResNet200 TopoFR Glint360K 97.84
PartialFC [79] ViT-B ArcFace WebFace42M 97.90
PartialFC [79] ResNet200 ArcFace WebFace42M 97.97
UniTSFace [66] ViT-L UniTSFace WebFace42M 97.99
TABLE VI: Performance on TinyFace [73] Dataset
@@
TransFace-L [147] ViT-S ArcFace MS1MV2 67.52 71.00
ARoFace [151] ResNet101 ArcFace MS1MV3 67.54 71.05
LGAF [150] ResNet101 ArcFace MS1MV2 68.35 71.59
ArcFace [20] ResNet101 ArcFace WebFace4M 71.11 74.38
AdaFace [22] ResNet101 AdaFace WebFace4M 72.02 74.52
AdaFace [22] ResNet101 AdaFace WebFace12M 72.29 74.97
REE [154] ResNet-50 ArcFace Native VLR 73.06 77.22
ARoFace [151] ResNet101 ArcFace WebFace4M 73.80 76.53
ARoFace [151] ResNet101 AdaFace WebFace4M 73.98 76.47
ARoFace [151] ResNet101 AdaFace WebFace12M 74.00 76.87
KP-RPE [39] ViT-B CosFace WebFace4M 75.48 78.30
KP-RPE [39] ViT-B ArcFace WebFace4M 75.62 78.57
KP-RPE [39] ViT-B AdaFace WebFace4M 75.80 78.49
TABLE VII: Performance on IJB-S [25] Dataset
Method Name Backbone Loss Function Training Data Rank-1 Rank-5
PFE [155] ResNet101 PFE MS1MV2 50.16 58.33
URL [149] ResNet101 URL MS1MV2 59.79 65.78
CurricularFace [21] ResNet101 CurricularFace MS1MV2 62.43 68.68
AdaFace [22] ResNet101 AdaFace MS1MV2 65.26 70.53
AdaFace [22] ViT AdaFace MS1MV3 65.95 71.64
AdaFace [22] ResNet101 AdaFace MS1MV3 67.12 72.67
KP-RPE [39] ViT AdaFace MS1MV3 67.62 73.25
ArcFace [20] ResNet101 ArcFace WebFace4M 69.26 74.31
AdaFace [22] ResNet101 AdaFace WebFace4M 70.42 75.29
ARoFace [151] ResNet101 ArcFace WebFace4M 70.96 75.54
AdaFace [22] ResNet101 AdaFace WebFace12M 71.35 76.24
AdaFace [22] ViT AdaFace WebFace4M 71.90 77.09
KP-RPE [39] ViT CosFace WebFace4M 72.22 77.67
ARoFace [151] ResNet101 AdaFace WebFace12M 72.28 77.93
KP-RPE [39] ViT AdaFace WebFace4M 72.78 78.20
KP-RPE [39] ViT ArcFace WebFace4M 73.04 78.62
@@ 5.1 Benchmark Evaluations
Robust evaluation of modern FR systems necessitates standardized benchmark datasets that reflect various real-world challenges. Prominent evaluation datasets extensively cited in recent literature include Labeled Faces in the Wild (LFW) [26], CFP-FP [27], CP-LFW [28], AgeDB [29], YouTube Faces (YTF) [156], TinyFace [73], and several iterations of the IARPA Janus Benchmark (IJB) series such as IJB-B [23], IJB-C [24], and IJB-S [25]. Each dataset addresses specific challenges inherent in FR scenarios. Below some datasets are described in detail.
Figure 14: Examples of faces across different datasets: IJB-S [25], TinyFace [73], IJB-C [24], and CFP-FP [27]. The top row shows gallery images, the middle row shows corresponding probe images, and the bottom row reports cosine similarity scores. Higher similarity scores (in green) indicate successful matches, while lower scores (in red) indicate mismatches. Features are extracted by KP-RPE [39] trained on WF4M [55].
@@
Labeled Faces in the Wild (LFW [26]) consists of over 13,000 facial images collected from the web, annotated with identity labels. The dataset includes multiple images for approximately 1,680 individuals of high-quality images. Main usage of this dataset is for the verification task.
@@
YouTube Faces (YTF [156]) specifically targets video-based unconstrained face recognition. The dataset comprises clips varying from 48 to 6,070 frames, with an average length of 181 frames and contains an average of 2 videos per subject, making it useful for assessing algorithms designed to handle real-world variability in videos.
@@
CFP-FP [27] evaluates the capability of algorithms to match frontal face images with their corresponding profile ones. It is particularly challenging due to large variations in facial orientation. The dataset is widely used to benchmark algorithms designed for pose-invariant face verification.
@@
TinyFace [73] is explicitly designed for low-resolution FR research at scale. It includes 169,403 naturally low-resolution images (average size 20x16 pixels) depicting 5,139 identities. Images in TinyFace are cropped from crowded scenes and span a diverse range of lighting, occlusion, backgrounds.
@@
IARPA Janus Benchmark-C (IJB-C) expands upon earlier series IJB-B [23], containing imagery and videos for 3,531 subjects, including 1,661 newly added identities. It comprises approximately 138,000 images and 11,000 videos. IJB-C serves as a challenging dataset for template-based recognition tasks. It contains significant variations in pose, illumination, and image quality [24]. Performance is often reported as TAR@FAR=threshold where the FAR threshold is selected based on the target operating point.
@@
IARPA Janus Surveillance Video Benchmark (IJB-S) targets surveillance-specific scenarios, featuring images and videos of 202 subjects collected at a Department of Defense facility. The galleries are comprised of high-quality upper torso images and the probes are videos captured by surveillance camera of varied altitude and range. It is suited for evaluating surveillance-oriented FR approaches [25].
@@
Collectively, these datasets represent comprehensive benchmarks that drive progress in addressing the nuanced challenges of modern face recognition technologies.
Figure 15: Evolution of SoTA performance in NIST FRVT 1:1 verification since January 2022. Each plot shows the cumulative minimum False Non-Match Rate (FNMR, lower is better) achieved by any submitted algorithm up to the corresponding date for the Visa, Mugshot, Visa Border, and Border datasets (the plot titles indicate gallery - probe in order). Performance is evaluated at a low False Match Rate (e.g., FMR=$10^{-6}$). The dashed red line indicates the performance level achieved by the WebFace42M entry (R200 [55] WF42M) for comparison.

### v3

@@ 4.5 Feature Fusion in Face Recognition
In template-based FR, multiple face images of the same identity—often captured under varying conditions of pose, illumination, resolution, and occlusion—must be fused into a single, compact representation to enable efficient and accurate comparison. This fusion scenario commonly arises in gallery settings, where multiple still images (or media) per subject must be aggregated into a unified template.
@@
A second, and increasingly popular, scenario involves video-based FR, where frames extracted from probe video sequences are fused into a single representation. This use case poses unique challenges, as it often requires online (on-the-fly) feature fusion to support real-time applications such as surveillance or mobile authentication. Despite differing temporal constraints, both still-image and video-based fusion share the core objective: to generate robust and compact representations that preserve discriminative identity cues.
@@
Feature fusion is a critical step in this process, as it determines how the information from diverse images of the same person is aggregated into a unified descriptor. Naive methods like average or max pooling treat all feature embeddings equally, which can dilute discriminative cues by giving equal importance to low-quality or redundant images. Effective feature fusion must not only compress but also intelligently filter, weight, and adapt to the content of the input set. The ability to generate order-invariant, and compact template representations directly impacts FR performance, especially in unconstrained or real-time scenarios.
@@
Early video-based FR approaches employed adaptive hidden Markov models to capture temporal dynamics and recognize entire video sequences [tu2009optimizing]. Over time, feature fusion in FR advanced from simple averaging to adaptive and context-aware neural aggregation. The Neural Aggregation Network (NAN) [yang2017neural] demonstrated the effectiveness of learning quality-aware attention weights for robust, order invariant face templates. Building on this, Multicolumn Networks [xie2018multicolumn] and C-FAN [gong2019video] introduced fine-grained quality analysis by modeling visual and contextual importance or by weighting individual feature channels. These developments significantly improved FR performance on challenging template-based benchmarks such as IJB-C [ijbc].
@@
Recent work emphasizes scalable and generalized feature fusion across diverse conditions. Methods like CAFace [kim2022cluster], CoNAN [jawade2023conan], and ProxyFusion [jawade2024proxyfusion] sustain performance even with templates containing many varied images. Practical approaches such as Norm Pooling [nanduri2024template] show that simple heuristics can be effective in multi domain settings, especially with limited training data. Overall, these innovations reflect a trend toward scalable, efficient strategies that handle long videos while maintaining robustness.
@@ 5 State of the art in Face Recognition
FR evaluation progressed from lab specific testing to standardized protocols [phillips1998feret, phillips2000feret, phillips2005overview]. Before deep learning, small proprietary datasets with fewer than 100 subjects and over 95% reported accuracy lacked generalizability [turk1991face, belhumeur1997eigenfaces, martinez2001pca]. The 1996 FERET program established large scale standardized evaluation, enabling systematic benchmarking and transparent comparison [phillips1998feret, phillips2000feret]. This shift to structured evaluation allowed consistent progress tracking and set stage for deep learning’s transformative impact on FR [krizhevsky2012imagenet, taigman2014deepface, schroff2015facenet, guo2019survey, wang2021deep, wang2022survey].
The FERET program [phillips1998feret] in the 1990s standardized FR benchmarks with 14,126 images of 1,199 individuals. A decade later, the FRGC [phillips2005overview] expanded this with 50,000 high resolution images and 3D scans, defining protocols that guided later NIST FRVT [grother2019face] series. Together, they formed the basis of modern FR benchmarks.
TABLE III: Performance on CFP-FP [cfpfp] Dataset
@@
GFace [zhao2025global] IResNet-50 GCE (LO) Casia-WebFace 97.44
CosFace [wang2018cosface] ResNet100 CosFace [wang2018cosface] MS1MV2 98.13
ArcFace [deng2019arcface] ResNet101 ArcFace MS1MV2 98.27
MV-Softmax [wang2020mis] ResNet100 MV-Softmax MS1MV2 98.28
CurricularFace [huang2020curricularface] ResNet101 CurricularFace MS1MV2 98.37
TransFace-B [dan2023transface] ResNet100 ArcFace MS1MV2 98.39
MagFace [meng2021magface] ResNet100 MagFace MS1MV2 98.46
AdaFace [kim2022adaface] ResNet101 AdaFace MS1MV2 98.49
CQA-Face [wang2022cqa] ResNet100 CQA-Face MS1MV2 98.49
UniFace [zhou2023uniface] ResNet100 UniFace [zhou2023uniface] MS1MV2 98.63
URL [shi2020towards] ResNet101 URL MS1MV2 98.64
LGAF [wang2025local] ResNet100 ArcFace MS1MV2 98.77
ArcFace [deng2019arcface] ResNet50 ArcFace Glint360K 98.77
ViT-S [dosovitskiy2020image] ViT-S ArcFace Glint360K 98.85
@@
TransFace-S [dan2023transface] ViT-S ArcFace Glint360K 98.91
AdaFace [kim2022adaface] ViT AdaFace WebFace4M 98.94
KP-RPE [kim2024keypoint] ViT AdaFace WebFace4M 99.01
ViT-B [dosovitskiy2020image] ViT-B ArcFace Glint360K 99.02
AdaFace [kim2022adaface] ResNet101 AdaFace MS1MV3 99.03
@@
AdaFace [kim2022adaface] ViT AdaFace MS1MV3 99.06
ArcFace [deng2019arcface] ResNet101 ArcFace WebFace4M 99.06
KP-RPE [kim2024keypoint] ViT ArcFace WebFace4M 99.09
@@
KP-RP [kim2024keypoint]E ViT AdaFace MS1MV3 99.11
GFace [zhao2025global] IResNet-100 GCE (LO) MS1MV3 99.12
@@
AdaFace [kim2022adaface] ResNet101 AdaFace WebFace4M 99.17
TransFace-B [dan2023transface] ViT-B ArcFace Glint360K 99.17
AdaFace [kim2022adaface] ResNet101 AdaFace WebFace12M 99.24
KP-RPE [kim2024keypoint] ViT AdaFace WebFace12M 99.30
TransFace-L [dan2023transface] ViT-L ArcFace Glint360K 99.32
TABLE IV: Performance on IJB-C [ijbc] Dataset
@@
ArcFace [deng2019arcface] ResNet101 ArcFace [deng2019arcface] MS1MV2 96.03
MagFace [meng2021magface] ResNet101 MagFace [meng2021magface] MS1MV2 95.81
@@
CurricularFace [huang2020curricularface] ResNet101 CurricularFace MS1MV2 96.10
@@
TransFace-S [dan2023transface] ViT-S ArcFace MS1MV2 96.45
TransFace-B [dan2023transface] ViT-B ArcFace MS1MV2 96.55
TransFace-L [dan2023transface] ViT-L ArcFace MS1MV2 96.59
@@
ARoFace [saadabadi2024aroface] ResNet101 ArcFace MS1MV2 96.66
ElasticFace [boutros2022elasticface] ResNet101 ElasticFace MS1MV2 96.65
TopoFR [dan2024topofr] ResNet101 TopoFR MS1MV2 96.95
GFace [zhao2025global] ResNet101 TopoFR MS1MV2 96.96
AdaFace [kim2022adaface] ResNet101 AdaFace MS1MV2 97.09
KP-RPE [kim2024keypoint] ViT-B CosFace WebFace4M 96.98
TopoFR [dan2024topofr] ResNet200 TopoFR MS1MV2 97.08
AdaFace [kim2022adaface] ViT-B AdaFace MS1MV3 97.10
KP-RPE [kim2024keypoint] ViT-B AdaFace WebFace4M 97.13
AdaFace [kim2022adaface] ViT-B AdaFace WebFace4M 97.14
KP-RPE [kim2024keypoint] ViT-B AdaFace MS1MV3 97.16
KP-RPE [kim2024keypoint] ViT-B ArcFace WebFace4M 97.21
PartialFC [an2021partial] ResNet101 ArcFace WebFace4M 97.22
CatFace [talemi2024catface] ResNet101 CatFace MS1MV2 97.43
AdaFace [kim2022adaface] ResNet101 AdaFace WebFace4M 97.39
ARoFace [saadabadi2024aroface] ResNet101 AdaFace WebFace4M 97.51
AdaFace [kim2022adaface] ResNet101 AdaFace WebFace12M 97.66
PartialFC [an2021partial] ResNet101 ArcFace WebFace12M 97.58
ARoFace [saadabadi2024aroface] ResNet101 AdaFace WebFace12M 97.60
TopoFR [dan2024topofr] ResNet101 TopoFR Glint360K 97.60
KP-RPE [kim2024keypoint] ViT-B AdaFace WebFace12M 97.82
PartialFC [an2021partial] ResNet101 ArcFace WebFace42M 97.82
TopoFR [dan2024topofr] ResNet200 TopoFR Glint360K 97.84
PartialFC [an2021partial] ViT-B ArcFace WebFace42M 97.90
PartialFC [an2021partial] ResNet200 ArcFace WebFace42M 97.97
UniTSFace [jia2023unitsface] ViT-L UniTSFace WebFace42M 97.99
TABLE V: Performance on IJB-S [ijbs] Dataset
Method Name Backbone Loss Function Training Data Rank-1 Rank-5
PFE [shi2019probabilistic] ResNet101 PFE MS1MV2 50.16 58.33
URL [shi2020towards] ResNet101 URL MS1MV2 59.79 65.78
CurricularFace [huang2020curricularface] ResNet101 CurricularFace MS1MV2 62.43 68.68
AdaFace [kim2022adaface] ResNet101 AdaFace MS1MV2 65.26 70.53
AdaFace [kim2022adaface] ViT AdaFace MS1MV3 65.95 71.64
AdaFace [kim2022adaface] ResNet101 AdaFace MS1MV3 67.12 72.67
KP-RPE [kim2024keypoint] ViT AdaFace MS1MV3 67.62 73.25
ArcFace [deng2019arcface] ResNet101 ArcFace WebFace4M 69.26 74.31
AdaFace [kim2022adaface] ResNet101 AdaFace WebFace4M 70.42 75.29
ARoFace [saadabadi2024aroface] ResNet101 ArcFace WebFace4M 70.96 75.54
AdaFace [kim2022adaface] ResNet101 AdaFace WebFace12M 71.35 76.24
AdaFace [kim2022adaface] ViT AdaFace WebFace4M 71.90 77.09
KP-RPE [kim2024keypoint] ViT CosFace WebFace4M 72.22 77.67
ARoFace [saadabadi2024aroface] ResNet101 AdaFace WebFace12M 72.28 77.93
KP-RPE [kim2024keypoint] ViT AdaFace WebFace4M 72.78 78.20
KP-RPE [kim2024keypoint] ViT ArcFace WebFace4M 73.04 78.62
TABLE VI: Performance on TinyFace [cheng2018low] Dataset
@@
TransFace-L [dan2023transface] ViT-S ArcFace MS1MV2 67.52 71.00
ARoFace [saadabadi2024aroface] ResNet101 ArcFace MS1MV3 67.54 71.05
LGAF [wang2025local] ResNet101 ArcFace MS1MV2 68.35 71.59
ArcFace [deng2019arcface] ResNet101 ArcFace WebFace4M 71.11 74.38
AdaFace [kim2022adaface] ResNet101 AdaFace WebFace4M 72.02 74.52
AdaFace [kim2022adaface] ResNet101 AdaFace WebFace12M 72.29 74.97
REE [chai2023recognizability] ResNet-50 ArcFace Native VLR 73.06 77.22
ARoFace [saadabadi2024aroface] ResNet101 ArcFace WebFace4M 73.80 76.53
ARoFace [saadabadi2024aroface] ResNet101 AdaFace WebFace4M 73.98 76.47
ARoFace [saadabadi2024aroface] ResNet101 AdaFace WebFace12M 74.00 76.87
KP-RPE [kim2024keypoint] ViT-B CosFace WebFace4M 75.48 78.30
KP-RPE [kim2024keypoint] ViT-B ArcFace WebFace4M 75.62 78.57
KP-RPE [kim2024keypoint] ViT-B AdaFace WebFace4M 75.80 78.49
@@ 5.1 Benchmark Evaluations
Robust evaluation of modern FR systems necessitates standardized benchmark datasets that reflect various real-world challenges. Prominent evaluation datasets extensively cited in recent literature include Labeled Faces in the Wild (LFW) [lfw], CFP-FP [cfpfp], CP-LFW [cplfw], AgeDB [agedb], YouTube Faces (YTF) [wolf2011face], TinyFace [cheng2018low], and several iterations of the IARPA Janus Benchmark (IJB) series such as IJB-B [ijbb], IJB-C [ijbc], and IJB-S [ijbs]. Each dataset addresses specific challenges inherent in FR scenarios. Below some datasets are described in detail.
@@
Labeled Faces in the Wild (LFW [lfw]) consists of over 13,000 facial images collected from the web, annotated with identity labels. The dataset includes multiple images for approximately 1,680 individuals of high-quality images. Main usage of this dataset is for the verification task.
@@
YouTube Faces (YTF [wolf2011face]) specifically targets video-based unconstrained FR. The dataset comprises clips varying from 48 to 6,070 frames, with an average length of 181 frames and contains an average of 2 videos per subject, making it useful for assessing algorithms designed to handle real-world variability in videos.
@@
CFP-FP [cfpfp] evaluates the capability of algorithms to match frontal face images with their corresponding profile ones. It is particularly challenging due to large variations in facial orientation. The dataset is widely used to benchmark algorithms designed for pose-invariant face verification.
@@
TinyFace [cheng2018low] is explicitly designed for low-resolution FR research at scale. It includes 169,403 naturally low-resolution images (average size 20x16 pixels) depicting 5,139 identities. Images in TinyFace are cropped from crowded scenes and span a diverse range of lighting, occlusion, backgrounds.
@@
IARPA Janus Benchmark-C (IJB-C) expands upon earlier series IJB-B [ijbb], containing imagery and videos for 3,531 subjects, including 1,661 newly added identities. It comprises approximately 138,000 images and 11,000 videos. IJB-C serves as a challenging dataset for template-based recognition tasks. It contains significant variations in pose, illumination, and image quality [ijbc]. Performance is often reported as TAR@FAR=threshold where the FAR threshold is selected based on the target operating point.
[…]
@@
IARPA Janus Surveillance Video Benchmark (IJB-S) targets surveillance-specific scenarios, featuring images and videos of 202 subjects collected at a Department of Defense facility. The galleries are comprised of high-quality upper torso images and the probes are videos captured by surveillance camera of varied altitude and range. It is suited for evaluating surveillance-oriented FR approaches [ijbs].
@@
Collectively, these datasets represent comprehensive benchmarks that drive progress in addressing the nuanced challenges of modern FR technologies.